Determine the nature of the solutions of the equation.
y^2=1/8y+3/5

Answer by stanbon(57387)

Determine the nature of the solutions of the equation.
y^2=(1/8)y+3/5
----
Multiply thru by 40 to get:
40y^2 = 5y + 24
---
40y^2-5y+24 = 0
----
Check the determinant:
b^2-4ac = 25-4*50*24 < 0
---
So It has two unequal complex roots.
============
Cheers,
Stan H.